OpenStudy (anonymous):

x(b-3)+(b-3) can someone explain to me how to factor these? I know its a simple question but its a bit hard for me

jimthompson5910 (jim_thompson5910):

You can factor out the common factor of b-3 to get x(b-3)+(b-3) (x+1)(b-3) So the answer is (x+1)(b-3)

hero (hero):

x(b-3)+(b-3) = x(b-3)+1(b-3) = (b-3)(b+1)

OpenStudy (anonymous):

Sure thing, pretend there is an invisible 1 after the plus sign since you have b-3 in both parenthesis, it can be factored out, so part of the answer is (b-3) Since you have an x and that invisible 1, the next part of the answer is (x+1) so your answer is (x+1)(b-3)

hero (hero):

You don't have to pretend the 1 is invisible, just add the 1 before the second binomial, then factor out (b-3)